HETAV PATEL

Grade 11,

Water drops are falling regularly at constant interval of 1s from a tap. If total time taken by a drop to strike is 4s,then distance between 2nd and 3rd drop when 1st drop just hits the ground is?

Water drops are falling regularly at constant interval of 1s from a tap. If total time taken by a drop to strike is 4s,then distance between 2nd and 3rd drop when 1st drop just hits the ground is?

Grade:12th pass

1 Answers

jitender
6 years ago
Time taken by 1st drop to reach groundS=ut+1/2 at^2 ( 2nd equation of motion)Now t= 4secu=0And a=g=10m/s^2S=height of tap=80mNow 2nd drop covers s2= 1/2 ×10×9=45mS3=1/2 ×10×4=20mBecause interval of dropping is 1 sec.S2-s3=45-20=25m